# Magnetocalorics of singlet ground state induced moment magnets

### v1

@@ III.1 Static field dependent susceptibility
An alternative approach to obtain $T_{m}(h)$ and the critical field $h_{c}(\xi_{s})$ defined by $T_{m}(h_{c})$=00 proceeds via the determination of collective static RPA susceptibilities. In the Ising-type singlet-singlet system only the operator $J_{z}=m_{s}S_{x}$ has matrix elements and leads to a finite static single-site susceptibility . For finite field and temperature it is given by [3, 15] (sublattice index $\lambda$ suppressed, $\tau,\tau^{\prime}=\pm$):
@@
$+$ $\beta\bigl{[}\sum_{\tau}|\langle\tau|S_{x}|\tau\rangle|^{2}-\langle S_{x}\rangle|^{2}\bigr{]}$
@@
$\{S_{x}\}_{\tau\tau^{\prime}}^{\lambda}=\frac{1}{2}\left(\begin{array}[]{cc}\sin 2\theta_{\lambda}&\cos 2\theta_{\lambda}\\ \cos 2\theta_{\lambda}&-\sin 2\theta_{\lambda}\\ \end{array}\right)$ (20)
@@
$\chi_{0\lambda}^{C}(T,h^{\prime})$ $=$ $(\frac{1}{2}m_{s})^{2}\frac{1}{T}(1-\frac{1}{(\hat{\Delta}_{T}^{\lambda})^{2}})(1-f_{s}^{\lambda 2})$ (21)
@@
Figure 5: Specific heat for FM with $\xi_{s}=1.2$ which is identical to AFM for $h=0$. However for finite field the FM order changes rapidly into a crossover (cf. Fig. 1(a)) and the sharp jump broadens and merges with the Schottky background for larger field. This type of field dependence should be expected for the induced FM PrI3 shown in Fig. 6 (only for zero field case).
Figure 6: Specific heat coefficient $\gamma=C_{V}/T$ of the the close to critical singlet-singlet induced ferromagnet at zero field. Dashed black line is the background Schottky anomaly and red line corresponds to control parameter $\xi_{s}=1.08$ close to the quantum critical $\xi^{c}_{s}$= 1. The jump at FM $T_{m}\equiv T_{C}$ appears at the maximum postion $T_{max}$ of the Schottky background $C^{0}_{V}/T$ (dashed black line). For comparison the magnetic specific heat of PrI3 (TC=7.5 K) adapted from Ref. 16 is shown by the red symbols (with residual T=0 $\gamma$-value subtracted and the maximum position and value scaled to coincide).
[…]
@@
$\text{AFM}:\hat{\chi}({\bf q})$ $=$ $\frac{1}{D({\bf q})}\left(\begin{array}[]{cc}\chi_{0A}&I({\bf q})\chi_{0A}\chi_{0B}\\ I({\bf q})\chi_{0B}\chi_{0A}&\chi_{0B}\\ \end{array}\right)$ (25)
$D({\bf q})$ $=$ $1-I({\bf q})^{2}({\bf q})\chi_{0A}\chi_{0B}$ (26)
@@
$\chi(0)=\frac{1}{2}\sum_{\lambda}\chi_{0\lambda}[1+I_{e}\chi_{0\lambda}]^{-1}$ (27)
As zero-field limiting cases we obtain for T= 0, Tm the following values [15] (in units of $[m_{s}^{2}/\Delta$]): For FM $\chi(0)=1/[2\xi_{s}(\xi_{s}^{2}-1)]$ and $\chi(T_{m})\rightarrow\infty$ while for AFM $\chi(0)=1/[2\xi_{s}(\xi_{s}^{2}+1)]$ and $\chi(T_{m})=1/(4\xi_{s})$. Note this is the ’longitudinal’ zz- type susceptibility (there is no transversal one in this Ising-type model) and it stays finite for T0absent0$\rightarrow$ 0→ 0. This is different from degenerate ground state magnets where it vanishes in that limit. For AFM the relative drop from the maximum at Tm may be expressed as
@@
Starting from the paramagnetic side of the AFM without sublattices we also may determine the transition temperature $T_{m}(h^{\prime})$ at finite field or AFM phase boundary from the condition that the AFM staggered susceptibility (q=Q) diverges, i.e. $I(({\bf Q})\chi_{0}(T_{m})=1$ with $I({\bf Q})=z|I_{0}|=I_{e}$ where ${\bf Q}$ is the AFM ordering vector. Using Eq. (21) this leads to the implicit solution for $T_{m}(h^{\prime})$ according to
$f_{s}(\hat{T}_{m})=\frac{\hat{\Delta}_{T_{m}}^{3}}{\xi_{s}}-\frac{1}{2}\frac{\hat{\Delta}_{T_{m}}}{\hat{T}_{m}}(\hat{\Delta}_{T_{m}}^{2}-1)\bigl{[}1-f_{s}^{2}(\hat{T}_{m})\bigr{]}$ (29)
with the normalized $\hat{T}_{m}=T_{m}(h^{\prime})/\Delta$ and paramagnetic level splitting $\hat{\Delta}_{T_{m}}(h^{\prime})=[1+(m_{s}h^{\prime}+2\xi_{s}\langle S_{x}\rangle)^{2}]^{\frac{1}{2}}$. The first and second terms on the r.h.s. of this equation result from vanVleck and pseudo Curie terms of the susceptibility, respectively. In zero field limit the latter vanishes and $\hat{\Delta}_{T_{m}}$=1, then the explicit solution of Eq. (16) is recovered
Furthermore the critical field for AFM order for a given control parameter $\xi_{s}>1$ defined by $T_{m}(h^{\prime}_{c})$= 0 can be obtained from Eq. (29) by setting $f_{s}(\hat{T}_{m}(h^{\prime}_{c}))$= 1 leading to the implicit equation
$h_{c}^{\prime}(\xi_{s})=\frac{1}{m_{s}}\bigl{[}(\xi_{s}^{\frac{2}{3}}-1)^{\frac{1}{2}}+2\xi_{s}\langle S_{x}\rangle_{0}\bigr{]}$ (30)
@@
$h^{\prime}_{c}=h_{c}(\delta)/\Delta$ $\simeq$ $\frac{\sqrt{2}}{m_{s}}(1+\frac{1}{\sqrt{3}})\sqrt{\delta}$ (31)
@@
Figure 7: Contour plots of FM (a,c) and AFM (b,d) specific heat $C_{V}(T,h)$ for two different control parameters $\xi_{s}$ with zero-field transition temperatures $T_{m}(\xi_{s}=1.1)=0.33$ and $T_{m}(\xi_{s}=1.2)=0.42$. In the FM case a transition occurs only for $h$= 0 and changes into a crossover with the Schottky type specific heat maximum at $T_{max}(h)$ shifting to higher temperature. In the AFM case the transition persists at all fields at $T_{m}(h)$ up to the critical fields $h^{\prime}_{c}(\xi_{s}=1.1)=0.53$ and $h^{\prime}_{c}(\xi_{s}=1.2)=0.77$ (cf. Fig. 2). The relative zero-field position of $T_{m}$ and background $T_{max}$ depends on $\xi_{s}$ (it coincides for $\xi_{s}=1.2$ in (d)) and they separate on approaching the critical field. Equidistant contour lines of $C_{V}(T,h)$ vary between 0-0.5 kB/site for $x_{s}=1.1$ and 0-0.75 kB/site for $\xi_{s}=1.2$
@@ IV Caloric properties and coefficients
With the basic properties of the model derived above we can now calculate the thermodynamic potentials as a prerequisite for obtaining the various thermodynamic quantities relevant for experimental investigation of singlet induced moment magnets. These potentials depend on the set of variables $(T,h,\varepsilon)$ where T is the temperature, h the external field and here $\varepsilon$ is the volume strain $\varepsilon$=$\frac{1}{3}$($\varepsilon_{xx}$+$\varepsilon_{yy}$+$\varepsilon_{zz}$) which enters through the magnetoelastic strain dependence $\Delta(\varepsilon)$ of the singlet-singlet splitting.
@@ IV.1 The thermodynamic potentials
$=\frac{1}{2}\sum_{\lambda}\sum_{n}p^{n}_{\lambda}\epsilon^{n}_{\lambda}+\frac{1}{2}\sigma m_{s}^{2}I_{e}\langle S_{x}\rangle_{\lambda}\langle S_{x}\rangle_{\bar{\lambda}}$ (32)
@@
$+\frac{1}{2}\sigma m_{s}^{2}I_{e}\langle S_{x}\rangle_{\lambda}\langle S_{x}\rangle_{\bar{\lambda}}$
@@
$U(T,h^{\prime})$ $=$ $-\frac{\Delta}{2}\bigl{\{}\hat{\Delta}_{T}f_{s}(T)-\frac{1}{2\xi_{s}}[(\hat{\Delta}_{T}^{2}-1)^{\frac{1}{2}}-m_{s}h^{\prime}]^{2}\bigr{\}}$
@@
$\langle S_{x}\rangle$ $=$ $\frac{\hat{\Delta}_{T}}{2}(m_{s}h^{\prime}+2\xi_{s}\langle S_{x}\rangle)\tanh\bigl{[}(\frac{\Delta}{2T})\hat{\Delta}_{T}\bigr{]}$ (34)
@@
$U(T,0)=-\frac{\Delta}{4}\bigl{[}\xi_{s}f_{s}^{2}(T)+\frac{1}{\xi_{s}}\bigr{]}$ (35)
@@
$S=\frac{U_{0}}{T}+lnZ=-\bigl{(}\frac{\Delta}{2T}\bigr{)}\sum_{\lambda}\hat{\Delta}^{\lambda}_{T}f_{s}^{\lambda}(T)+\frac{1}{2}\sum_{\lambda}Z_{\lambda}$ (36)
@@
$S(T,h^{\prime})$ $=$ $\frac{1}{2}\sum_{\lambda}\ln\Bigl{\{}2\cosh\bigl{[}\bigl{(}\frac{\Delta}{2T}\bigr{)}\hat{\Delta}_{T}^{\lambda}\bigr{]}$ (37)
$-\bigl{(}\frac{\Delta}{2T}\bigr{)}\hat{\Delta}^{\lambda}_{T}\tanh\bigr{[}\bigl{(}\frac{\Delta}{2T}\bigr{)}\hat{\Delta}_{T}^{\lambda}\bigr{]}\Bigl{\}}$
@@
$F(T,h^{\prime})=$ $-T\frac{1}{2}\sum_{\lambda}\ln\Bigl{\{}2\cosh\bigl{(}\frac{\Delta}{2T}\bigr{)}\hat{\Delta}_{T}^{\lambda}\Bigr{\}}$ (38)
$+\frac{1}{2}\sigma m_{s}^{2}I_{e}\langle S_{x}\rangle_{\lambda}\langle S_{x}\rangle_{\bar{\lambda}}$
@@
Figure 9: Magnetocaloric coefficients (adiabatic cooling rate with magnetic field) for FM and AFM cases with $\xi_{s}=1.2$. For FM a monotonic increase for lowering the field is observed $(T<T_{m})$ which diverges on reaching $T_{m}$ from below. For temperatures $T>T_{m}$ (dashed) it will be suppressed to zero in the low field limit. For the AF case a sign change and downward peak at a fields with $T_{m}(h)=T$ occurs and $\Gamma_{mc}$ approaches zero at low fields. For temperatures above $T_{m}$ (dashed) a featureless sub- linear field dependence occurs.
IV.2 Specific heat, magneto- and elastocaloric coefficients and thermal expansion
@@
$C_{V}(T,h^{\prime})=\bigl{(}\frac{\partial U(T,h^{\prime})}{\partial T}\bigr{)}_{\varepsilon}$ (39)
@@
Further interesting and important experimentally accessible quantities are the magneto- and elasto- (or baro-) caloric coefficients, i.e. the adiabatic (S=const) cooling rates under rapid field or hydrostatic pressure change, respectively [17, 18, 19]. Microscopically the former is due to the field dependence of non-diagonal elements in the Hamiltonian of Eq. (4) while the latter is caused by the volume strain dependence of diagonal terms (i.e. the CEF splitting). The adiabatic magnetocaloric coefficient is given by
$\Gamma_{mc}(T,H)=\bigl{(}\frac{\partial T}{\partial H}\bigr{)}_{S}=-\frac{T}{C_{V}}\bigl{(}\frac{\partial M_{x}}{\partial T}\bigr{)}_{H}$ (40)
with the homogeneous magnetization $M_{x}$ = $m_{s}(g_{J}\mu_{B})M_{x}^{0}$ = $(g_{eff}\mu_{B})M_{x}^{0}$ and $M_{x}^{0}$ defined in Eq. (11). Likewise we have for the adiabatic barocaloric coefficient at a constant field H
$\Gamma_{bc}(T,H)=\bigl{(}\frac{\partial T}{\partial p}\bigr{)}_{S}=\frac{T}{C_{p}}\bigl{(}\frac{\partial\varepsilon}{\partial T}\bigr{)}_{p}$ (41)
where $p=-c_{B}\varepsilon$ is the pressure and the last factor in this equation is the thermal expansion coefficient with $c_{B}$ denoting the inverse bulk modulus or fully symmetric elastic constant.
$\beta(T,H)=\bigl{(}\frac{\partial\varepsilon}{\partial T}\bigr{)}_{p};\;\;\epsilon(T)=-\frac{1}{c_{B}}\bigl{(}\frac{\partial F}{\partial\varepsilon}\bigr{)}_{T}$ (42)
which is also an important accessible quantity in CEF systems [20] and can be obtained from the free energy as indicated. We assume that the strain coupling appears predominantly through the volume strain dependence of the CEF splitting $\Delta(\epsilon)$. Then a dimensionless CEF Grüneisen parameter $\Omega=-(\partial ln\Delta(\epsilon)/\partial\epsilon)$ (which may be positive or negative) characterizes the strength of this coupling. In rare earth compounds this parameter is usually of the order of $|\Omega|\approx$ 10. [21, 20]. Then we may write for the strain dependent energy scale $\Delta(\epsilon)=\Delta(1-\Omega\epsilon)$ and control parameter $\xi_{s}(\epsilon)=\xi_{s}(1+\Omega\epsilon)$ and this leads to the strain dependence of the free energy in Eq. (38) which determines thermal expansion, magnetostriction and barocaloric coefficients given in the equations above.
Typical examples of the magneto- and barocaloric quantities for FM and AFM case are presented in Figs. 9,10 and discussed in Sec. VII.
@@
V The case of PrI3: a near-critical singlet induced ferromagnet

### v2

@@ III.1 Static field dependent susceptibility
An alternative approach to obtain $T_{m}(h)$ and the critical field $h_{c}(\xi_{s})$ defined by $T_{m}(h_{c})$=0 proceeds via the determination of collective static RPA susceptibilities. In the Ising-type singlet-singlet system only the operator $J_{z}=m_{s}S_{x}$ has matrix elements and leads to a finite static single-site susceptibility . For finite field and temperature it is given by [10, 3] (sublattice index $\lambda$ suppressed, $\tau,\tau^{\prime}=\pm$):
@@
$+$ $\beta\bigl{[}\sum_{\tau}|\langle\tau|S_{x}|\tau\rangle|^{2}-\langle S_{x}\rangle|^{2}\bigr{]},$
@@
$\{S_{x}\}_{\tau\tau^{\prime}}^{\lambda}=\frac{1}{2}\left(\begin{array}[]{cc}\sin 2\theta_{\lambda}&\cos 2\theta_{\lambda}\\ \cos 2\theta_{\lambda}&-\sin 2\theta_{\lambda}\\ \end{array}\right).$ (20)
@@
$\chi_{0\lambda}^{C}(T,h^{\prime})$ $=$ $(\frac{1}{2}m_{s})^{2}\frac{1}{T}(1-\frac{1}{(\hat{\Delta}_{T}^{\lambda})^{2}})(1-f_{s}^{\lambda 2}).$ (21)
@@
Figure 5: Specific heat coefficient $\gamma=C_{V}/T$ of the the close to critical singlet-singlet induced ferromagnet at zero field. Dashed black line is the background Schottky anomaly and red line corresponds to control parameter $\xi_{s}=1.08$ close to the quantum critical $\xi^{c}_{s}$= 1. The jump at FM $T_{m}\equiv T_{C}$ appears at the maximum postion $T_{max}$ of the Schottky background $C^{0}_{V}/T$ (dashed black line). For comparison the magnetic specific heat of PrIr3 (TC=7.5 K) adapted from Ref. 4 is shown by the red symbols (with residual T=0 $\gamma$-value subtracted and the maximum position and value scaled to coincide).
$\text{FM}:\;\;$ $\chi({\bf q})=\chi_{0}[1-I({\bf q})\chi_{0}]^{-1},$ (22)
@@
$\hskip-11.38092pt\text{AFM}:\hat{\chi}({\bf q})$ $=$ $\frac{1}{D({\bf q})}\left(\begin{array}[]{cc}\chi_{0A}&I({\bf q})\chi_{0A}\chi_{0B}\\ I({\bf q})\chi_{0B}\chi_{0A}&\chi_{0B}\\ \end{array}\right)$ (25)
$D({\bf q})$ $=$ $1-I({\bf q})^{2}\chi_{0A}\chi_{0B}.$ (26)
@@
$\chi(0)=\frac{1}{2}\sum_{\lambda}\chi_{0\lambda}[1+I_{e}\chi_{0\lambda}]^{-1}.$ (27)
As zero-field limiting cases we obtain for T= 0, Tm the following values [3] (in units of $m_{s}^{2}/\Delta$]): For FM $\chi(0)=1/[2\xi_{s}(\xi_{s}^{2}-1)]$ and $\chi(T_{m})\rightarrow\infty$ while for AFM $\chi(0)=1/[2\xi_{s}(\xi_{s}^{2}+1)]$ and $\chi(T_{m})=1/(4\xi_{s})$. Note this is the ’longitudinal’ zz- type susceptibility shown in Fig. 3 (there is no transversal one in this Ising-type model) and it stays finite for $\rightarrow 0. T$his is different from degenerate ground state magnets where it vanishes in that limit. For AFM the relative drop from the maximum at Tm may be expressed as
@@
Starting from the paramagnetic side of the AFM without sublattices we also may determine the transition temperature $T_{m}(h^{\prime})$ at finite field or AFM phase boundary from the condition that the AFM staggered susceptibility (q=Q) diverges, i.e. $I({\bf Q})\chi_{0}(T_{m})=1$ with $I({\bf Q})=z|I_{0}|=I_{e}$ where $\bf Q$} is the AFM ordering vector. Using Eq. (21) this leads to the implicit solution for $T_{m}(h^{\prime})$ according to
$f_{s}(\hat{T}_{m})=\frac{\hat{\Delta}_{T_{m}}^{3}}{\xi_{s}}-\frac{1}{2}\frac{\hat{\Delta}_{T_{m}}}{\hat{T}_{m}}(\hat{\Delta}_{T_{m}}^{2}-1)\bigl{[}1-f_{s}^{2}(\hat{T}_{m})\bigr{]},$ (29)
with the normalized $\hat{T}_{m}=T_{m}(h^{\prime})/\Delta$ and paramagnetic level splitting $\hat{\Delta}_{T_{m}}(h^{\prime})=[1+(m_{s}h^{\prime}+2\xi_{s}\langle S_{x}\rangle)^{2}]^{\frac{1}{2}}$. The first and second terms on the r.h.s. of this equation result from vanVleck and pseudo Curie terms of the susceptibility, respectively. In zero field limit the latter vanishes and $\hat{\Delta}_{T_{m}}$=1, then the explicit solution of Eq. (16) is recovered.
Furthermore the critical field for AFM order for a given control parameter $\xi_{s}>1$ is defined by the condition $T_{m}(h^{\prime}_{c})$= 0. At zero temperature the population difference in Eq.(13) reaches its maximum, meaning $f_{s}(\hat{T}_{m}(h^{\prime}_{c}))$= 1. Inserting this into Eq. (29) leads to the implicit equation for the critical field according to
$h_{c}^{\prime}(\xi_{s})=\frac{1}{m_{s}}\bigl{[}(\xi_{s}^{\frac{2}{3}}-1)^{\frac{1}{2}}+2\xi_{s}\langle S_{x}\rangle_{0}\bigr{]},$ (30)
@@
$h^{\prime}_{c}=h_{c}(\delta)/\Delta$ $\simeq$ $\frac{\sqrt{2}}{m_{s}}(1+\frac{1}{\sqrt{3}})\sqrt{\delta}.$ (31)
@@
Figure 6: Contour plots of FM (a,c) and AFM (b,d) specific heat $C_{V}(T,h)$ for two different control parameters $\xi_{s}$ with zero-field transition temperatures $T_{m}(\xi_{s}=1.1)=0.33$ and $T_{m}(\xi_{s}=1.2)=0.42$. In the FM case a transition occurs only for hh= 0 and changes into a crossover with the Schottky type specific heat maximum at $T_{max}(h)$ shifting to higher temperature. In the AFM case the transition persists at all fields at $T_{m}(h)$ up to the critical fields $h^{\prime}_{c}(\xi_{s}=1.1)=0.53$ and $h^{\prime}_{c}(\xi_{s}=1.2)=0.77$ (cf. Fig. 2). The relative zero-field position of $T_{m}$ and background $T_{max}$ depends on $\xi_{s}$ (it coincides for $\xi_{s}=1.2$ in (d)) and they separate on approaching the critical field. Equidistant contour lines of $C_{V}(T,h)$ vary between 0-0.5 kB/site for $x_{s}=1.1$ and 0-0.75 kB/site for $\xi_{s}=1.2$ (see Fig. 4).
@@ IV Caloric properties and coefficients
With the basic properties of the model derived above we can now calculate the thermodynamic potentials as a prerequisite for obtaining the various thermodynamic quantities relevant for experimental investigation of singlet induced moment magnets. These potentials depend on the set of variables $(T,h,\varepsilon)$ where T is the temperature, h the external field and $\varepsilon$ is the volume strain $\varepsilon$$\frac{1}{3}$($\varepsilon_{xx}$$\varepsilon_{yy}$$\varepsilon_{zz}$) which enters through the magnetoelastic strain dependence $\Delta(\varepsilon)$ of the singlet-singlet splitting.
@@ IV.1 The thermodynamic potentials
$=\frac{1}{2}\sum_{\lambda}\sum_{n}p^{n}_{\lambda}\epsilon^{n}_{\lambda}+\frac{1}{2}\sigma m_{s}^{2}I_{e}\langle S_{x}\rangle_{\lambda}\langle S_{x}\rangle_{\bar{\lambda}},$ (32)
@@
$+\frac{1}{2}\sigma m_{s}^{2}I_{e}\langle S_{x}\rangle_{\lambda}\langle S_{x}\rangle_{\bar{\lambda}}.$
@@
$\hskip-8.5359ptU(T,h^{\prime})$ $=$ $-\frac{\Delta}{2}\bigl{\{}\hat{\Delta}_{T}f_{s}(T)-\frac{1}{2\xi_{s}}[(\hat{\Delta}_{T}^{2}-1)^{\frac{1}{2}}-m_{s}h^{\prime}]^{2}\bigr{\}}$
@@
$\langle S_{x}\rangle$ $=$ $\frac{\hat{\Delta}_{T}}{2}(m_{s}h^{\prime}+2\xi_{s}\langle S_{x}\rangle)\tanh\bigl{[}(\frac{\Delta}{2T})\hat{\Delta}_{T}\bigr{]}.$ (34)
@@
$U(T,0)=-\frac{\Delta}{4}\bigl{[}\xi_{s}f_{s}^{2}(T)+\frac{1}{\xi_{s}}\bigr{]},$ (35)
@@
$\hskip-11.38092ptS=\frac{U_{0}}{T}+lnZ=-\bigl{(}\frac{\Delta}{2T}\bigr{)}\sum_{\lambda}\hat{\Delta}^{\lambda}_{T}f_{s}^{\lambda}(T)+\frac{1}{2}\sum_{\lambda}Z_{\lambda},$ (36)
@@
$S(T,h^{\prime})$ $=$ $\frac{1}{2}\sum_{\lambda}\Bigl{\{}\ln\Bigl{[}2\cosh\bigl{[}\bigl{(}\frac{\Delta}{2T}\bigr{)}\hat{\Delta}_{T}^{\lambda}\bigr{]}\Bigr{]}$ (37)
$-\bigl{(}\frac{\Delta}{2T}\bigr{)}\hat{\Delta}^{\lambda}_{T}\tanh\bigr{[}\bigl{(}\frac{\Delta}{2T}\bigr{)}\hat{\Delta}_{T}^{\lambda}\bigr{]}\Bigl{\}},$
@@
$F(T,h^{\prime})=$ $-T\frac{1}{2}\sum_{\lambda}\ln\Bigl{[}2\cosh\bigl{(}\frac{\Delta}{2T}\bigr{)}\hat{\Delta}_{T}^{\lambda}\Bigr{]}$ (38)
$+\frac{1}{2}\sigma m_{s}^{2}I_{e}\langle S_{x}\rangle_{\lambda}\langle S_{x}\rangle_{\bar{\lambda}}.$
@@
[…]
IV.2 Specific heat, magneto- and barocaloric coefficients and thermal expansion
@@
$C_{V}(T,h^{\prime})=\bigl{(}\frac{\partial U(T,h^{\prime})}{\partial T}\bigr{)}_{\varepsilon},$ (39)
@@
Further interesting and important experimentally accessible quantities are the magneto- and elasto- (or baro-) caloric coefficients, i.e. the adiabatic (S=const) cooling rates under rapid field or hydrostatic pressure change, respectively [11, 12, 13]. These effects have also technical applications which are not in the focus in this work and we refer to Refs.14, 15 for a deeper discussion of this important issue. Microscopically the magneto- and barocaloric coefficient are due to the field dependence of non-diagonal elements in the Hamiltonian of Eq. (4) or the volume strain dependence of diagonal terms (i.e. the CEF splitting), respectively. The adiabatic magnetocaloric coefficient is given by
$\Gamma_{mc}(T,H)=\bigl{(}\frac{\partial T}{\partial H}\bigr{)}_{S}=-\frac{T}{C_{V}}\bigl{(}\frac{\partial M_{x}}{\partial T}\bigr{)}_{H},$ (40)
with the homogeneous magnetisation $M_{x}$ = $m_{s}(g_{J}\mu_{B})M_{x}^{0}$ = $(g_{eff}\mu_{B})M_{x}^{0}$ and $M_{x}^{0}$ defined in Eq. (11). Likewise we have for the adiabatic barocaloric coefficient at a constant field H
$\Gamma_{bc}(T,H)=\bigl{(}\frac{\partial T}{\partial p}\bigr{)}_{S}=\frac{T}{C_{p}}\bigl{(}\frac{\partial\varepsilon}{\partial T}\bigr{)}_{p},$ (41)
where $p=-c_{B}\varepsilon$ is the pressure and the last factor in the above equation is the thermal expansion coefficient $\beta$ with $c_{B}$ denoting the inverse bulk modulus or fully symmetric elastic constant.
$\beta(T,H)=\bigl{(}\frac{\partial\varepsilon}{\partial T}\bigr{)}_{p};\;\;\epsilon(T)=-\frac{1}{c_{B}}\bigl{(}\frac{\partial F}{\partial\varepsilon}\bigr{)}_{T}.$ (42)
These are also an important accessible quantities in CEF systems [16] and can be obtained from the free energy as indicated. We assume that the strain coupling appears predominantly through the volume strain dependence of the CEF splitting $\Delta(\epsilon)$. Then a dimensionless CEF Grüneisen parameter $\Omega=-(\partial ln\Delta(\epsilon)/\partial\epsilon)$ (which may be positive or negative) characterises the strength of this coupling. In rare earth compounds this parameter is usually of the order of $|\Omega|\approx$ 10. [17, 16]. Then we may write for the strain dependent energy scale $\Delta(\epsilon)=\Delta(1-\Omega\epsilon)$ and control parameter $\xi_{s}(\epsilon)=\xi_{s}(1+\Omega\epsilon)$ and this leads to the strain dependence of the free energy in Eq. (38) which determines thermal expansion, magnetostriction and barocaloric coefficients given in the equations above.
Typical examples of the magneto- and barocaloric quantities for FM and AFM case are discussed in Sec. VII.
@@
V Material examples for singlet induced moment magnets
Now we give a brief overview of materials where the theory of field dependent properties in singlet ground state induced quantum magnets developed here may be relevant and then apply it to a most recent example of the near critical ferromagnet PrIr3.
In this overview we also include materials where the first excited CEF state is not a singlet but may also be a doublet (for four- and sixfold uniaxial point groups) or a triplet (for tetrahedral and cubic point groups only). It was demonstrated in Ref. 3 that their principal thermodyamic properties are similar to the singlet-singlet case. As shown in the following section the induced moment magnetism may also be preceded in the paramagnetic phase by a strong temperature dependence and a softening of a critical magnetic exciton mode to a varying degree at the ordering wave vector. A considerable number of materials have been accumulated that qualify as singlet ground state induced moment system. Frequently these compounds contain non-Kramers 4f or 5f ions like (J=4) Pr3Tl[18, 19], PrSb [20], PrCu2[21], PrNi[22], Pr5Ge4 [23], Pr2CuO4 [24] and elemental Pr- metal under pressure [10, 25, 26, 8]. Singlet ground state induced AFM order has also been found in Tb and Tm compounds (both with J=6) TbSb [27] and its Y- substituted alloy [9] as well as Tb3Ga5O12 [28] and KTmSe3 [29]. Recently a 4f singlet-singlet system PrMgAl11O19 [30] has been discovered with Pr located on a triangular sublattice. This raises the interesting theoretical question how the induced moment mechanism interferes with geometric frustration. Indeed this compound does not order magnetically down to $0.4$ K.
Furthermore candidates with 5f electrons for induced magnetism are UGa2 [31] and UPd2Al3 [32, 33, 34] and URu2Si2[35] (further references cited therein), and also Fe-substituted [36]. However, it should be noted that the degree of itineracy of 5f states is often higher and the localised picture may not be fully applicable. When considered within the localised 5f scenario URu2Si2[35, 36] and also the (Kramers ion) compound YbRu2Ge2 [37] show that the induced order mechanism is not restricted to magnetism but also possible for multipolar [38, 39] and (ferro-)quadrupolar [40, 41] order, respectively.
Finally another aspect of f- electron materials with CEF singlet ground state has been discovered: For suitable 2D lattice structure with a basis like honeycomb or kagome the magnetic excitons may develop a nontrivial topological character with non-vanishing Chern number which requires the presence of magnetic excitonic edge modes in the paramagnetic state [42]. An example is the 2D honeycomb lattice with a two-atom basis with site symmetry $C_{3v}$. which may be realised as a planar structure within a 3D lattice. This structure is relevant for various f-electron compounds like Na2PrO3 [43], TmNi3Al9[44] and recently a new class of promising 4f (RE =Tm (J=6),Ho (J=8)) honeycomb materials BaRE2(SiO4)6 was found [45]. All of these compounds have integer total angular momentum $J$ and thus will exhibit a CEF singlet ground state in this site symmetry.
V.1 The case of PrIr3 : a new near-critical singlet induced ferromagnet